# Weighted Ensemble Models Are Strong Continual Learners

### v1

@@ 3.2 Continual Model Averaging
Temporarily disregarding the typical memory limitations of CIL, a practical approach through model ensembling is proposed as follows: individual models are trained for each task sequentially, starting with the pre-trained model, and the model is saved after each task. Suppose we have trained on a sequence of $T$ tasks, then this method will result in $T$ distinct neural networks, each with its own set of parameters, denoted as $\theta_{1},\ldots,\theta_{T}$. The objective is to create a composite neural network with parameters $\theta$ ensuring effective performance across all the tasks.
@@
We model the posterior over the composite parameters $\theta$ conditioned on the task-specific parameters $\theta_{t}$ as $p(\theta|\theta_{t})$. The parameters are assumed to be distributed according to an isotropic Gaussian distribution $\theta\sim\mathcal{N}({\theta}_{t},I)$, where $I$ is the identity matrix [75, 38]. By maximizing the average of the task-specific log-likelihoods of these posterior distributions for all $T$ models, we obtain the following optimization problem:
$\theta^{*}=\arg\max_{\theta}\frac{1}{T}\sum_{t=1}^{T}\log p(\theta|\theta_{t}).$ (1)
@@
The solution to this optimization problem has closed-form and is the simple average of the model parameters [38]. With a slight abuse in notation, where we use the sum operator to denote the element-wise summation across sets, it can be written:
$\theta^{*}=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}.$ (2)
Below, we detail how we adapt this model-averaging method to the constraints of CIL where the task data arrives sequentially, and without storing previous data.
First, one of the goals of our approach is to prevent the number of parameters from growing linearly with $T$. Therefore, we shift from simultaneous averaging all models $\theta_{1},\ldots,\theta_{t}$ to the iterative computation of the average $\theta^{*}$:
$\theta^{*}_{t}=\lambda\theta_{t}+(1-\lambda)\theta^{*}_{t-1},$ (3)
where $\lambda\!\in\![0,1]$. We add model-level weighting $\lambda$ as an additional hyper-parameter to set the relative importance of each model. In terms of memory, during each training phase, the storage overhead of our approach compared to naive sequential fine-tuning over all the tasks is limited to the size of a single model. However, when transitioning to subsequent tasks, only $\theta^{*}_{t}$ requires storage. To initialize this recursion, we start at $t\!=\!1$ with $\theta^{*}_{1}$ which is set to the model parameters obtained at the end of the first task. Moreover, the rationale for Eq. 3 is reinforced by the findings of Mirzadeh et al. [41]. This work indicates that, given two models each trained on a separate task, a model proficient in both tasks typically resides at the linear interpolation of their respective parameter spaces.
@@
Secondly, initiating the training of each task from the same pre-trained model $\theta_{0}$ can result in convergence to distinct regions within the parameter space [16, 53]. This case is illustrated in Fig. 2 (a). However, if the learned parameters are too distant, the Gaussian assumption of the posterior distribution is no longer valid, and averaging the network could lead to high-loss regions and poor performance of the aggregated model. Therefore, for each task $t$, we initiate finetuning from $\theta^{*}_{t-1}$ rather than the initial pre-trained model $\theta_{0}$. This change limits the risk of reaching distant parameter regions.
Third, we observe that the uniform averaging of the models may still lead to a region that is far from every $\theta_{t}$ as illustrated in Fig. 2 (b). Therefore, we propose to perform the non-uniform averaging, giving higher importance to the latest tasks as shown in Fig. 2 (c). This is obtained by using a constant weight parameter $\lambda\!=\!0.5$. The motivation for giving more importance to the latest tasks is that early models are trained only on the first few tasks, while the last model has encoded the knowledge from both old and recent tasks via sequential fine-tuning. In this way, the final model $\theta^{*}_{T}$ is unlikely to fall in the region with the high value of the loss since it remains close to the model from the last task.
@@
Handling Classifier Parameters. In each novel task $t$, there are unique parameters (specifically, new head parameters associated with new classes) not present in the preceding models that are subject to averaging. To accommodate this, we restrict the averaging process (as in Eq. 3) exclusively to the parameters that are common across models (this includes both backbone parameters and the shared portions of head parameters), while excluding the new head weights (pertaining to new classes) from the averaging.
Algorithm 1 CoFiMA
Inputs: Training dataset $\mathcal{D}_{t}$ for task $t=1,...,T$; network $M_{\theta}(\cdot)$ with pre-trained parameters $\theta_{0}^{*}$;
Initialization: Random Initialization of the classifiers.
1:# sequential tasks.
2:for task $t=1,...,T$ do
3:     - Train $M_{\theta_{t}}$ with cross-entropy loss on $\mathcal{D}_{t}$.
4:     - Calculate Fisher Diagonal information $F_{t}$ for $\theta_{t}$.
5:     - Calculate new weights $\theta^{*}_{t}$ according to Eq. 7.
6:     - Calculate Fisher Diagonal information $F_{t}^{*}$ for $\theta^{*}_{t}$.
7:end for
@@ 3.3 Continual Fisher-weighted Model Averaging (CoFiMA)
Uniform weight-averaging operates under the implicit assumption that all weights of the model have the same importance for the training task $t$. This assumption could potentially compromise model performance since different parameters can have various impacts on the network output [28]. To enhance the averaging process, we propose a more refined model averaging for CL. Specifically, we employ the Fisher information matrix [59, 10], which encapsulates the quantity of information that observed data $x$ provides about the network parameters $\theta_{t}$. The Fisher information matrix $F_{\theta_{t}}$ of a neural network $p_{\theta_{t}}(y|x)$ trained to predict an output $y$ from input data $x$ is is computed as in [3]:
$F_{\theta_{t}}=\mathbb{E}_{x{\sim}\mathcal{D}_{t}}\big{\{}\mathbb{E}_{y\sim p(y|x,\theta_{t})}[$ $\nabla_{\theta_{t}}\log p(y|x,\theta_{t})$
$\nabla_{\theta_{t}}\log p(y|x,\theta_{t})^{T}]\big{\}}.$ (4)
Given the computational cost of storing the full Fisher matrix, this study adopts the diagonal of the Fisher matrix for practicality [47, 59, 28]. The computation of the diagonal Fisher is feasible within the same order of complexity as the standard back-propagation training process, as it necessitates only a single gradient computation per data point. Through Monte Carlo sampling [59, 60], the diagonal Fisher matrix can be estimated as follows:
$\hat{F}_{\theta_{t}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{y\sim p(y|x_{i},\theta_{t})}\left\{[\nabla_{\theta_{t}}\log p(y|x_{i},\theta_{t})]^{2}\right\},$ (5)
where the expectation over $y$ is estimated from the data samples $N$. The estimation of this Fisher information matrix $F_{\theta}$ is based on the assumption that the statistical properties of the data are locally similar around the parameter values [58, 35]. This assumption aligns with the Gaussian assumption previously utilized in deriving CoMA. However, as discussed in the work of Chizat et al. [6], the use of PTMs limits the risk of facing significant changes in parameter distributions which would violate the locality assumption [10].
We now consider that the posterior of each task model $p(\theta|\theta_{t})$ is defined as $\theta\sim\mathcal{N}({\theta}_{t},F_{t}^{-1})$. As in Eq. 2, maximization of the average of the task-specific log-likelihoods brings us to the following closed-form solution:
$\theta^{*}=\frac{\sum_{t=1}^{T}F_{t}\theta_{t}}{\sum_{t=1}^{T}F_{t}}$ (6)
@@
$\theta_{t}^{*}=\frac{\lambda F_{t}\theta_{t}+(1-\lambda)F^{*}_{t-1}\theta^{*}_{t-1}}{\lambda F_{t}+(1-\lambda)F^{*}_{t-1}},$ (7)
where $F^{*}_{t-1}$ is the Fisher matrix estimated with $\theta^{*}_{t-1}$ on the data $\mathcal{D}_{t-1}$. In terms of memory requirements, this solution requires the storage of only the model parameters $\theta^{*}_{t}$ and the Fisher matrix $\theta_{t}^{*}$ in between two consecutive tasks. Similarly to CoMA, the recursion is initialized at the end of the first task with $\theta^{*}_{1}\!=\!\theta_{1}$ and estimating $F^{*}_{1}$ with $\theta_{1}$ on $\mathcal{D}_{1}$.
Method Memory Split CIFAR-100 Split ImageNet-R Split CUB-200 Split Cars-196
Joint-Training - 93.22±$\pm 0.16$ 79.60±$\pm 0.87$ 88.00±$\pm 0.34$ 80.31±$\pm 0.13$
GDumb [49] 81.92±$\pm 0.15$ 89.46±$\pm 0.94$ 24.23±$\pm 0.35$ 43.48±$\pm 0.49$ 61.80±$\pm 0.77$ 79.76±$\pm 0.18$ 25.20±$\pm 0.84$ 49.48±$\pm 0.74$
DER++ [4] 84.50±$\pm 1.67$ 91.49±$\pm 0.61$ 67.75±$\pm 0.93$ 78.13±$\pm 1.14$ 77.42±$\pm 0.71$ 87.61±$\pm 0.09$ 60.41±$\pm 1.76$ $\pm 0.57$
BiC [77] 88.45±$\pm 0.57$ 93.37±$\pm 0.32$ 64.89±$\pm 0.80$ 73.66±$\pm 1.61$ 81.91±$\pm 2.59$ 89.29±$\pm 1.57$ 63.10±$\pm 5.71$ 73.75±$\pm 2.37$
L2P [73] $\checkmark$ 82.76±$\pm 1.17$ 88.48±$\pm 0.83$ 66.49±$\pm 0.40$ 72.83±$\pm 0.56$ 62.21±$\pm 1.92$ 73.83±$\pm 1.67$ 38.18±$\pm 2.33$ 51.79±$\pm 4.19$
DualPrompt [72] $\checkmark$ 85.56±$\pm 0.33$ 90.33±$\pm 0.33$ 68.50±$\pm 0.52$ 72.59±$\pm 0.24$ 66.00±$\pm 0.57$ 77.92±$\pm 0.50$ 40.14±$\pm 2.36$ 56.74±$\pm 1.78$
EWC [28] $\checkmark$ 89.30±$\pm 0.23$ 92.31±$\pm 1.66$ 70.27±$\pm 1.99$ 76.27±$\pm 2.13$ 68.32±$\pm 2.64$ 79.95±$\pm 2.28$ 52.50±$\pm 3.18$ 64.01±$\pm 3.25$
LwF [34] $\checkmark$ 87.99±$\pm 0.05$ 92.13±$\pm 1.16$ 67.29±$\pm 1.67$ 74.47±$\pm 1.48$ 69.75±$\pm 1.37$ 80.45±$\pm 2.08$ 49.94±$\pm 3.24$ 63.28±$\pm 1.11$
Prototype-classifier 60.29±$\pm 0.00$ 69.18±$\pm 0.00$ 38.45±$\pm 0.00$ 45.59±$\pm 0.00$ 80.66±$\pm 0.00$ 88.95±$\pm 0.00$ 28.58±$\pm 0.00$ 39.83±$\pm 0.00$
Seq FT $\checkmark$ 88.86±$\pm 0.83$ 92.01±$\pm 1.71$ 71.80±$\pm 1.45$ 76.84±$\pm 1.26$ 68.07±$\pm 1.09$ 79.04±$\pm 1.69$ 49.74±$\pm 1.25$ 62.83±$\pm 2.16$
SLCA [82] $\checkmark$ 91.53±$\pm 0.28$ 94.09±$\pm 0.87$ 77.00±$\pm 0.33$ 81.17±$\pm 0.64$ 84.71±$\pm 0.40$ 90.94±$\pm 0.68$ 67.73±$\pm 0.85$ 76.93±$\pm 1.21$
CoFiMA (Ours) 92.16±$\pm 0.24$ 94.66±$\pm 0.94$ 79.48±$\pm 0.26$ 83.28±$\pm 0.56$ 87.02±$\pm 0.56$ 92.03±$\pm 0.69$ 73.46±$\pm 0.64$ 79.35±$\pm 0.96$
Table 1: Experimental results for continual learning on Split CIFAR-100, Split ImageNet-R, Split CUB-200 and Split Cars-196 using ViT-B/16 [11] supervised pre-training on ImageNet-21K [54].
Method Memory Split CIFAR-100 Split ImageNet-R Split CUB-200 Split Cars-196
Joint-Training - 89.11±$\pm 0.06$ 72.80±$\pm 0.23$ 79.55±$\pm 0.04$ 74.52±$\pm 0.09$
GDumb [49] 69.72±$\pm 0.20$ 80.95±$\pm 1.19$ 28.24±$\pm 0.58$ 43.64±$\pm 1.05$ 45.29±$\pm 0.97$ 66.86±$\pm 0.63$ 20.95±$\pm 0.42$ 45.40±$\pm 0.66$
DER++ [4] 63.64±$\pm 1.30$ 79.55±$\pm 0.87$ 53.11±$\pm 0.44$ 65.10±$\pm 0.91$ 61.47±$\pm 0.32$ 77.15±$\pm 0.61$ 50.64±$\pm 0.70$ 67.64±$\pm 0.45$
BiC [77] 80.57±$\pm 0.86$ 89.39±$\pm 0.33$ 57.36±$\pm 2.68$ 68.07±$\pm 0.22$ 74.39±$\pm 1.12$ 82.13±$\pm 0.33$ 65.57±$\pm 0.93$ 73.95±$\pm 0.29$
EWC [28] $\checkmark$ 81.62±$\pm 0.34$ 87.56±$\pm 0.97$ 64.50±$\pm 0.36$ 70.37±$\pm 0.41$ 61.36±$\pm 1.43$ 72.84±$\pm 2.18$ 53.16±$\pm 1.45$ 63.61±$\pm 1.06$
LwF [34] $\checkmark$ 77.94±$\pm 1.00$ 86.90±$\pm 0.90$ 60.74±$\pm 0.30$ 68.55±$\pm 0.65$ 61.66±$\pm 1.95$ 73.90±$\pm 1.91$ 52.45±$\pm 0.48$ 63.87±$\pm 0.31$
Seq FT $\checkmark$ 81.47±$\pm 0.55$ 87.55±$\pm 0.95$ 64.43±$\pm 0.44$ 70.48±$\pm 0.54$ 61.67±$\pm 1.37$ 73.25±$\pm 1.83$ 52.91±$\pm 1.61$ 63.32±$\pm 1.31$
SLCA [82] $\checkmark$ 85.27±$\pm 0.08$ 89.51±$\pm 1.04$ 68.07±$\pm 0.21$ 73.04±$\pm 0.56$ 73.01±$\pm 0.16$ 82.13±$\pm 0.34$ 66.04±$\pm 0.08$ 72.59±$\pm 0.04$
CoFiMA (Ours) 86.59±$\pm 0.47$ 89.66±$\pm 0.53$ 68.15±$\pm 0.31$ 73.66±$\pm 0.78$ 75.86±$\pm 0.18$ 83.37±$\pm 0.16$ 63.93±$\pm 0.16$ 70.83±$\pm 0.83$
Table 2: Experimental results on Split CIFAR-100, Split ImageNet-R, Split CUB-200 and Split Cars-196 using MoCo v3 [5] self-supervised pre-training on ImageNet-1K.

### v3

@@ 3.2 Continual Model Averaging (CoMA)
Temporarily disregarding the typical memory limitations of CIL, a practical approach through model ensembling is proposed as follows: individual models are trained for each task sequentially, starting from the pre-trained model, and the model is saved after each task. Suppose we have trained on a sequence of $T$ tasks: this method will result in $T$ distinct networks, each with its own set of parameters, denoted as $\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{T}$. The objective is to create a composite neural network with parameters $\boldsymbol{\theta}$ ensuring good performance across all the tasks.
@@
We model the posterior over the composite parameters $\boldsymbol{\theta}$, conditioned on the task-specific parameters $\boldsymbol{\theta}_{t}$, denoted by $p(\boldsymbol{\theta}|\boldsymbol{\theta}_{t})$, as an isotropic Gaussian distribution $\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{\theta}_{t},\mbox{\bf I})$, where I is the identity matrix [80, 37]. We treat the models $\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{T}$ as independent observations of the composite model $\boldsymbol{\theta}$ and maximizing the log-likelihoods of the posterior distribution of $\boldsymbol{\theta}$ over all $T$ the tasks leads to the following optimization problem:
$\boldsymbol{\theta}^{*}=\arg\max_{\boldsymbol{\theta}}\frac{1}{T}\sum_{t=1}^{T}\log p(\boldsymbol{\theta}|\boldsymbol{\theta}_{t}).$ (1)
@@
The solution to this optimization problem has closed-form and is the simple average of the model parameters [39, 37]. With a slight abuse in notation, where we use the sum operator to denote the element-wise summation across sets, it can be written:
$\boldsymbol{\theta}^{*}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\theta}_{t}.$ (2)
In addition to the support of the well-established likelihood maximization framework, averaging models as in Eq. (2) gains further validation from the insights provided by Mirzadeh et al. [41]. Their research shows that when two models are each trained on distinct tasks, a model adept in both tasks often exists within the linear interpolation of their parameter spaces. Below, we detail how we adapt this model-averaging method to the constraints of CIL where the task data arrives sequentially, and without storing previous data.
@@
First, initiating the training of each task from the same pre-trained model $\boldsymbol{\theta}_{0}$ can result in convergence to distinct regions within the parameter space [14, 53]. This case is illustrated in Fig. 2(a). If the learned parameters are too distant, the Gaussian assumption of the posterior distribution is no longer valid, and averaging the network could lead to high-loss regions and poor performance of the aggregated model (see Sec. 4.3). Therefore, for each task $t$, we initiate finetuning from $\boldsymbol{\theta}^{*}_{t-1}$ rather than the initial pre-trained model $\boldsymbol{\theta}_{0}$. This change limits the risk of reaching distant parameter regions.
Second, one of the goals of our approach is to prevent the number of models from growing linearly with $t$. Therefore, we shift from simultaneous averaging all models $\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{T}$ to the iterative computation of the average $\boldsymbol{\theta}^{*}_{t}$ at each task $t$:
$\boldsymbol{\theta}^{*}_{t}=\lambda\boldsymbol{\theta}_{t}+(1-\lambda)\boldsymbol{\theta}^{*}_{t-1},$ (3)
where $\lambda\!\in\![0,1]$. To initialize this recursion, we start at $t\!=\!1$ with $\boldsymbol{\theta}^{*}_{1}$ which is set to the model parameters obtained at the end of the first task. Note that, when $\lambda=\frac{1}{t}$, Eq. (3) is strictly equivalent to the average in Eq. (2) (see supplementary material), aligning with the maximum-likelihood solution and adhering to our initial hypothesis of isotropic Gaussian posteriors. However, our experiments highlight that such a parameter choice might result in suboptimal performances and that assigning variable weights to each model proves advantageous. This phenomenon might be attributed to the possibility of encountering a region with a high loss value in between all the task-specific models (see Fig. 2(b)). Therefore, we propose to perform a non-uniform averaging giving higher importance to the latest tasks as shown in Fig. 2(c). This is obtained by using a constant weight parameter $\lambda$. The motivation for giving more importance to the latest tasks is that early models are trained only on the first few tasks, while the last model has encoded the knowledge from both old and recent tasks via sequential fine-tuning (see supplementary material). In this way, it is expected that a better trade-off would exist along this trajectory, and the two endpoints $\boldsymbol{\theta}_{t-1}$ and $\boldsymbol{\theta}_{t}$ are smoothly connected without significant loss barrier or performance drop along the path [14].
In terms of memory, during each training phase, the storage overhead of our approach compared to naive sequential fine-tuning over all the tasks is limited to the size of a single model. However, when transitioning to subsequent tasks, only $\boldsymbol{\theta}^{*}_{t}$ requires storage.
@@
Handling Classifier Parameters. In each novel task $t$, there are unique parameters (specifically, new head parameters associated with new classes) not present in the preceding models that are subject to averaging. To accommodate this, we restrict the averaging process (as in Eq. (3)) exclusively to the parameters that are common across models (this includes both backbone parameters and the shared portions of head parameters) while excluding the new head weights (pertaining to new classes) from the averaging.
@@ 3.3 Continual Fisher-weighted Model Averaging (CoFiMA)
Uniform weight-averaging operates under the implicit assumption that all parameters of the model have the same importance for the training task $t$. This assumption could potentially compromise model performance since different parameters can have various impacts on the network output [26, 37]. To enhance the averaging process, we propose a more refined model averaging for CIL. Specifically, we employ the Fisher information matrix [62, 9], which encapsulates the quantity of information that observed data x provides about the network parameters $\boldsymbol{\theta}_{t}$. The Fisher information matrix $F_{\boldsymbol{\theta}_{t}}$ of a neural network $p_{\boldsymbol{\theta}_{t}}(\boldsymbol{y}|\mbox{\bf x})$ trained to predict an output $\boldsymbol{y}$ from input data x is computed as in [3]:
$\mbox{\bf F}_{\boldsymbol{\theta}_{t}}=\mathbb{E}_{\mbox{\bf x}{\sim}\mathcal{D}_{t}}\big{[}\mathbb{E}_{\boldsymbol{y}\sim p(\boldsymbol{y}|\mbox{\bf x},\boldsymbol{\theta}_{t})}[$ $\nabla_{\boldsymbol{\theta}_{t}}\log p(\boldsymbol{y}|\mbox{\bf x},\boldsymbol{\theta}_{t})\nabla_{\boldsymbol{\theta}_{t}}\log p(\boldsymbol{y}|\mbox{\bf x},\boldsymbol{\theta}_{t})^{T}]\big{]}.$ (4)
Given the computational cost of storing the full Fisher matrix, this study adopts the diagonal of the Fisher information matrix for practicality [48, 62, 26]. The computation of the diagonal Fisher is feasible within the same order of complexity as the standard back-propagation training process, as it necessitates only a single gradient computation per data point, corresponding to an extra epoch at the end of each task training. Through Monte-Carlo sampling over x [62, 63], the diagonal Fisher matrix can be estimated as follows:
$\hat{F}_{\boldsymbol{\theta}_{t}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\boldsymbol{y}\sim p(\boldsymbol{y}|\mbox{\bf x}_{i},\boldsymbol{\theta}_{t})}\left[[\nabla_{\boldsymbol{\theta}_{t}}\log p(\boldsymbol{y}|\mbox{\bf x}_{i},\boldsymbol{\theta}_{t})]^{2}\right],$ (5)
where the expectation over $\boldsymbol{y}$ is estimated from the data samples $N$. The estimation of this Fisher information matrix $F_{\boldsymbol{\theta}}$ is based on the assumption that the statistical properties of the data are locally similar around the parameter values [61, 34]. This assumption aligns with the Gaussian assumption previously utilized in deriving CoMA. Furthermore, as discussed in the work of Chizat et al. [6], the use of PTMs limits the risk of facing significant changes in parameter distributions which would violate the locality assumption [9].
We consider that the posterior of each task model $p(\boldsymbol{\theta}|\boldsymbol{\theta}_{t})$ is defined as ${}{\boldsymbol{\theta}\!\sim\!\mathcal{N}({\boldsymbol{\theta}}_{t},\mbox{\bf F}_{t}^{-1})}$. As in Eq. (2), maximization of the average of the task-specific log-likelihoods brings us to the following closed-form solution:
$\boldsymbol{\theta}^{*}=\frac{\sum_{t=1}^{T}\mbox{\bf F}_{t}\boldsymbol{\theta}_{t}}{\sum_{t=1}^{T}\mbox{\bf F}_{t}},$ (6)
@@
$\boldsymbol{\theta}_{t}^{*}=\frac{\lambda\mbox{\bf F}_{t}\boldsymbol{\theta}_{t}+(1-\lambda)\mbox{\bf F}_{t-1}\boldsymbol{\theta}^{*}_{t-1}}{\lambda\mbox{\bf F}_{t}+(1-\lambda)\mbox{\bf F}_{t-1}},$ (7)
where $F_{t-1}$ is the Fisher matrix estimated with $\boldsymbol{\theta}^{*}_{t-1}$ on the data $\mathcal{D}_{t-1}$. In terms of memory requirements, this solution requires the storage of only the model parameters $\boldsymbol{\theta}^{*}_{t}$ and the Fisher matrix $\mbox{\bf F}_{t-1}$ in between two consecutive tasks $t$ and $t\!-\!1$. Similarly to CoMA, the recursion is initialized at the end of the first task with $\boldsymbol{\theta}^{*}_{1}\!=\!\boldsymbol{\theta}_{1}$ and estimating $\mbox{\bf F}_{1}$ with $\boldsymbol{\theta}_{1}$ on $\mathcal{D}_{1}$.